The Absolute Zero Co. just issued a dividend of $2.60 per share on its common stock The company is expected to maintain a constant 5.4 percent growth rate in its dividends indefinitely. 


If the stock sells for $52 a share, what is the company's cost of equity? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 


Cost of equity = _______ 
Answer #1

This question requires application of constant growth dividend discount model according to which:

$$ \begin{gathered} P_{0}=\frac{D i v_{1}}{r-g} \\ P_{0}=\text { Price of Stock } \\ \text { Div }_{1}=\text { Estimated Dividends for Next Period } \\ r=\text { Required Rate of Return } \\ g=\text { Growth Rate } \end{gathered} $$

$$ \begin{aligned} &\operatorname{Div} 1=\operatorname{Div} 0 *(1+g) \\ &\text { Div } 1=\$ 2.60 *(1+5.4 \%)=\$ 2.7404 \\ &52=\frac{2.7404}{r-0.054} \\ &r-0.054=0.0527 \end{aligned} $$

Cost of Equity, \(r=10.67 \%\)